2. Select a distinctive positive integer a with a > 10 that is not a perfect cube a) Use a third degree Taylor Polynomial tolet a = 35

Please show work!
Answer #1

2. Select a distinctive positive integer a with a > 10 that is not a perfect Use a third degree Taylor Polynomial to approxim3 3 No uw 3 23- 2. 3 27- 2.3 thu 2, 3 l 2. 2 3 5 8 = 3.271849556 ( 35.02514792 802 16 C. we cam see +hohe IS 1-353 = 7.83246018 × 10-4 3.27185.. ans